OpenStudy (anonymous):

linear equation using POINT-PLOTTING, INTERCEPTS & SLOPE-INTERCEPT METHOD GIVEN: 2x-y=4

OpenStudy (unklerhaukus):

can you re-arrange that equation so it is y= ...

OpenStudy (anonymous):

im sorry but that's the only given.. 2x-y=4

OpenStudy (unklerhaukus):

you could add y to both sides then take 4 away from both sides

OpenStudy (unklerhaukus):

what do you get ...

OpenStudy (anonymous):

y=-2x+4 is that right?

OpenStudy (unklerhaukus):

almost, check you signs -/+

OpenStudy (anonymous):

yah! that my answerr y=-2x+4 is that right?

OpenStudy (unklerhaukus):

2x-y=4 adding y to both sides 2x-y+y=4+y 2x = 4+y taking 4 away from both sides 2x -4 = 4-4+y =

OpenStudy (anonymous):

why adding both sides?? all i know is to subtract it? is that right?

OpenStudy (unklerhaukus):

the other way is 2x -y = 4 taking away 2x form both sides 2x-2x-y = 4 -2x -y = 4 -2x now multiply both sides by minus one